# Coupler Position Optimization and Channel Estimation for Flexible Coupler Aided Multiuser Communication

### v1

@@ IV Channel Estimation for Flexible Coupler Systems
As shown in Fig. 2, we propose a channel estimation protocol for FC systems. Specifically, the channel training length $T_{c}$ (in time slots) is divided into $V$ pilot blocks indexed by $v\in\{1,\ldots,V\}$. At pilot block $v$, FC antenna $m$ sets the positions of its associated $N$ couplers to $\mathbf{p}_{m}^{[v]}$. We assume that all user channels remain constant over $T_{c}=V\tau$, where $\tau$ denotes the number of time slots in each pilot block. Therefore, the coupler position vector $\mathbf{p}_{m}^{[v]}$ remains unchanged during the $\tau$ time slots of the $v$-th block, and it varies across different blocks. Each user transmits a known pilot sequence $\mathbf{S}]_{k,:}\in\mathbb{C}^{1\times\tau}$ to the BS, and the same pilot is repeated over the $V$ blocks. Let pilot matrix $\mathbf{S}\in\mathbb{C}^{K\times\tau}$ satisfy $\mathbf{S}\mathbf{S}^{\mathsf{H}}=\tau\mathbf{I}_{K}$.
Based on the proposed protocol, we first develop a pilot-assisted centralized channel estimation algorithm with high estimation accuracy and then present a pilot-assisted distributed channel estimation algorithm for FC systems.
@@ IV-A Centralized Channel Estimation
denote the local channel from user $k$ to the $m$-th FC antenna under coupler position $\mathbf{p}_{m}^{[v]}$, where $\mathbf{h}_{\mathrm{C},k}(\mathbf{p}_{m}^{[v]})=\sum_{\ell=1}^{L_{k}}\alpha_{k,\ell}\mathbf{a}_{\mathrm{C},k,\ell}(\mathbf{p}_{m}^{[v]})\in\mathbb{C}^{N\times 1}$ collects the $N$ coupler to user-$k$ channels associated with $\mathbf{p}_{m}^{[v]}$. Under the proposed pilot pattern, the received signal at FC antenna $m$ is given by
@@
where $\mathbf{H}_{m}(\mathbf{p}_{m}^{[v]})=[\mathbf{h}_{m,1}(\mathbf{p}_{m}^{[v]}),\ldots,\mathbf{h}_{m,K}(\mathbf{p}_{m}^{[v]})]\in\mathbb{C}^{(N+1)\times K}$ denotes the channel matrix between the $m$-th FC antenna and the users, and $\mathbf{n}_{m}^{[v]}\in\mathbb{C}^{1\times\tau}$ is the received Gaussian noise vector.
@@
In the centralized channel estimation scheme, for each pilot block $v$, every FC antenna sends its received signal $\mathbf{y}_{m}^{[v]}$ to the CPU. The CPU stacks the received pilots across the $M$ antennas as
@@
Moreover, according to (10), the effective channel from user $k$ to FC antenna $m$ at signal block $v$ can be written as
@@
The main steps of the proposed centralized channel estimation procedure are summarized in Algorithm 2. The computational complexity of Algorithm 2 is dominated by the OMP-based sparse recovery and scales as $\mathcal{O}\!\big(\sum_{k=1}^{K}L_{k}\,MV\,G\big)$, while constructing the dictionary incurs an additional $\mathcal{O}(MVGN)$ operations. Therefore, the overall complexity is $\mathcal{O}\!\big(\sum_{k=1}^{K}L_{k}\,MV\,G+MVGN\big)$. In terms of communication overhead, Algorithm 2 requires each FC antenna to upload its length-$\tau$ received pilot vectors $\{\mathbf{y}_{m}^{[v]}\}_{v=1}^{V}$ to the CPU over $V$ pilot blocks, resulting in a communication overhead on the order of $\mathcal{O}(MV\tau)$. The centralized channel estimation scheme provides high estimation accuracy since the CPU jointly processes the pilot-correlated observations stacked over all FC antennas and all pilot blocks.
Algorithm 2 Centralized channel estimation for FC system
@@
3:  Each FC antenna $m$ applies $\mathbf{p}_{m}^{[v]}$ and receives $\mathbf{y}_{m}^{[v]}$ in (73).
4:  Each FC antenna $m$ uploads $\mathbf{y}_{m}^{[v]}$ to the CPU.
@@ IV-B Distributed Channel Estimation
The centralized channel estimation approach suffers from prohibitively high computational complexity, processing cost, and latency at the CPU, especially when the BS is equipped with an extremely large FC array with a large number of active antennas $M$ and a vast number of coupler candidate positions. We aim to design a distributed channel estimation scheme for the FC array that achieves accuracy close to the centralized baseline, while reducing the CPU computational burden.
@@ IV-B1 Processing at local nodes
Similar to the centralized channel estimation step, we discretize the azimuth domain into $G\gg L_{k}$ grid points $\{\phi_{1},\ldots,\phi_{G}\}$ and define the local dictionary at FC antenna $m$ as
@@
where $\mathbf{x}_{m,k}\in\mathbb{C}^{G\times 1}$ is sparse and its support encodes the azimuth angles of user $k$ on the grid. The support is common across $m$ since these angles are global parameters shared by all FC antennas. Moreover, $\mathbf{n}_{m,k}\!\sim\!\mathcal{CN}(\mathbf{0},\sigma_{m,\mathrm{eff}}^{2}\mathbf{I}_{V})$ denotes the effective noise vector after pilot correlation and stacking over the $V$ pilot blocks, with $\sigma_{m,\mathrm{eff}}^{2}=\sigma_{m}^{2}/\tau$.
To convey support information with low overhead, FC antenna $m$ computes a matched-filter proxy $\mathbf{c}_{m,k}=\mathbf{A}_{m}^{\mathsf{H}}\widehat{\mathbf{y}}_{m,k}$ and its normalized energy $\rho_{m,k}(j)=\frac{|[\mathbf{c}_{m,k}]_{j}|^{2}}{\|\mathbf{a}_{m}(\phi_{j})\|_{2}^{2}+\epsilon_{\mathrm{n}}}$, $j=1,\ldots,G$, where $\epsilon_{\mathrm{n}}>0$ avoids numerical issues when $\|\mathbf{a}_{m}(\phi_{j})\|_{2}^{2}$ is small. It then selects a candidate support set via noise-calibrated hard thresholding [6]:
@@
where $\eta>0$ controls the pruning aggressiveness. Finally, FC antenna $m$ uploads only the index-value pairs $\mathcal{U}_{m,k}=\{(j,\rho_{m,k}(j))\}_{j\in\mathcal{I}_{m,k}}$ to the CPU. Since the azimuth domain is sparse, the uploaded list is typically low-dimensional, i.e., $|\mathcal{I}_{m,k}|\ll G$ in typical multipath channels.
@@ IV-B2 Processing at central node
Since azimuth angles $\{\phi_{k,\ell}\}_{\ell=1}^{L_{k}}$ are path-level geometric parameters and thus common to all FC antennas, they are global parameters shared across different FC antennas. Local index set $\{\mathcal{I}_{m,k}\}$ provides complementary evidence of this common support on the grid. Accordingly, the CPU forms an aggregated grid score for each $j\in\{1,\ldots,G\}$ as
@@ IV-B3 Post-processing at local nodes
The CPU broadcasts the estimated support set, $\widehat{\Omega}_{k}$, to all FC antennas. Given $\widehat{\Omega}_{k}$, FC antenna $m$ restricts its local dictionary to the columns indexed by $\widehat{\Omega}_{k}$, namely,
@@
where $\bm{\alpha}_{k}\in\mathbb{C}^{L_{k}\times 1}$ collects the path gains aligned with the ordering of indices in $\widehat{\Omega}_{k}$. The reduced model allows each FC antenna to contribute to estimating the path gains using only a few coefficients, instead of processing the full $G$-dimensional sparse vector.
Algorithm 3 Distributed channel estimation for FC systems
@@
4:  For each FC antenna $m$ in parallel, apply $\mathbf{p}_{m}^{[v]}$ and receive $\mathbf{y}_{m}^{[v]}$.
5:  For each FC antenna $m$ in parallel, compute $\widehat{g}_{m,k}^{[v]}$ for all $k$ via (87).
@@
7:for each FC antenna $m$ in parallel and each user $k$ do
@@
18:for each FC antenna $m$ in parallel and each user $k$ do
@@
To avoid sending the $V$-dimensional observation $\widehat{\mathbf{y}}_{m,k}$ from each FC antenna, FC antenna $m$ computes and forwards the following sufficient statistics for LS estimation to the CPU,
@@ IV-B4 Post-processing at central node
After receiving LS sufficient statistics $\{\mathbf{R}_{m,k},\mathbf{q}_{m,k}\}_{m=1}^{M}$ from all FC antennas, the CPU aggregates them as
@@
After obtaining $\{\widehat{\Omega}_{k},\widehat{\bm{\alpha}}_{k}\}_{k=1}^{K}$ from the CPU, FC antenna $m$ reconstructs its effective scalar channel under any coupler position $\mathbf{p}_{m}$ as in (86). The reconstructed channel state information (CSI) is utilized for the subsequent coupler position optimization and digital precoding procedures described in Section III.
@@ V Simulation Results
Figure 3: Channel power gain (dB) between the FC array and the user with $N=1$, $K=1$, $M=5$, $P=30$ dBm, and $L_{k}=15$.
In this section, numerical results are presented to evaluate the performance of the proposed distributed FC system design. The coupler movement region associated with each active antenna is modeled as a square with side length $A$. In the simulations, both the active antennas and the passive couplers are modeled as thin straight wire dipoles. The mutual impedance matrix is computed from closed-form expressions according to [1]. The active antennas are placed with an inter element spacing of $d_{y}=2.2\lambda$. Each coupler is allowed to move within a square region of side length $A=2\lambda$, subject to a minimum element spacing constraint of $d_{\min}=0.15\lambda$. The carrier frequency is set to $f_{c}=7$ GHz. The load-impedance matrix $\mathbf{X}$ is diagonal with identical entries $0.05+j50~\Omega$.
Figure 4: Achievable rate versus maximum transmit power for different schemes with $N=2$, $K=3$, $M=15$, $P=30$ dBm, and $L_{k}=15$.
@@ V-A Performance of Coupler Position Optimization
To evaluate the performance of coupler position optimization, the proposed FC scheme is compared with the following baseline schemes.
@@
  • Fixed-position coupler: The couplers are uniformly placed within each movement region, and the load impedance matrix $\mathbf{X}$ has constant diagonal entries of $0.05+j50~\Omega$.
  • Parasitic antenna: Closed-form optimization of load impedance matrix $\mathbf{X}$ [9] is used to maximize the sum rate while keeping the coupler positions fixed at the same positions as in the fixed-position coupler case.
  • Fully active array: All $MN+M$ ports are configured as active antennas and uniformly placed at fixed positions within each movement region.
@@
Fig. 4 plots the achievable sum rate versus maximum transmit power $P_{\max}$ for the proposed FC array and several baseline schemes. As $P_{\max}$ increases, all schemes achieve higher rates due to the improved effective signal-to-noise ratio (SNR). It is observed that the proposed FC array closely approaches the fully active array with $MN+M$ RF chains while using only $M$ active antennas. This gain is achieved because coupler position optimization provides substantial mechanical beamforming and fading mitigation gains in rich multipath channels while maintaining a limited RF chain cost. In contrast, the parasitic antenna scheme optimizes only the load impedances with fixed coupler locations and therefore cannot fully exploit the spatial degree of freedom (DoF) offered by position reconfiguration. Moreover, the fixed position coupler and the active antenna array exhibit lower rates due to their fixed geometries and the lack of a reconfigurable aperture.
Figure 5: Achievable rate versus number of users with $N=2$, $M=10$, $P=30$ dBm, and $L_{k}=15$.
In Fig. 5, we plot the achievable sum rate versus the number of users $K$ for the proposed FC array and benchmark schemes. As $K$ increases, the sum rate improves for all schemes due to the increased spatial multiplexing gain, while the growth gradually slows down since stronger multiuser interference must be mitigated under a fixed transmit power budget. Compared with the parasitic antenna and fixed position coupler schemes, the proposed design consistently achieves a higher sum rate. This performance advantage becomes more pronounced at moderate and large values of $K$, where MMSE based multiuser precoding is more sensitive to the conditioning of the effective channel matrix and thus benefits more from geometry reconfiguration, whereas the benchmark schemes with fixed geometries cannot adapt to user induced interference coupling.

### v2

@@ IV Channel Estimation for Flexible Coupler Systems
As shown in Fig. 2, we propose a channel estimation protocol for FCA systems. Specifically, the channel training length $T_{c}$ (in time slots) is divided into $V$ pilot blocks indexed by $v\in\{1,\ldots,V\}$. At pilot block $v$, FCA $m$ sets the positions of its associated $N$ couplers to $\mathbf{p}_{m}^{[v]}$. We assume that all user channels remain constant over $T_{c}=V\tau$, where $\tau$ denotes the number of time slots in each pilot block. Therefore, the coupler position vector $\mathbf{p}_{m}^{[v]}$ remains unchanged during the $\tau$ time slots of the $v$-th block, and it varies across different blocks. Each user transmits a known pilot sequence $\mathbf{S}]_{k,:}\in\mathbb{C}^{1\times\tau}$ to the BS, and the same pilot is repeated over the $V$ blocks. Let pilot matrix $\mathbf{S}\in\mathbb{C}^{K\times\tau}$ satisfy $\mathbf{S}\mathbf{S}^{\mathsf{H}}=\tau\mathbf{I}_{K}$.
Based on the proposed protocol, we first develop a pilot-assisted centralized channel estimation algorithm with high estimation accuracy and then present a pilot-assisted distributed channel estimation algorithm for FCA systems.
@@ IV-A Centralized Channel Estimation
denote the local channel from user $k$ to the $m$-th FCA under coupler position $\mathbf{p}_{m}^{[v]}$, where $\mathbf{h}_{\mathrm{C},k}(\mathbf{p}_{m}^{[v]})=\sum_{\ell=1}^{L_{k}}\alpha_{k,\ell}\mathbf{a}_{\mathrm{C},k,\ell}(\mathbf{p}_{m}^{[v]})\in\mathbb{C}^{N\times 1}$ collects the $N$ coupler to user-$k$ channels associated with $\mathbf{p}_{m}^{[v]}$. Under the proposed pilot pattern, the received signal at FCA $m$ is given by
@@
where $\mathbf{H}_{m}(\mathbf{p}_{m}^{[v]})=[\mathbf{h}_{m,1}(\mathbf{p}_{m}^{[v]}),\ldots,\mathbf{h}_{m,K}(\mathbf{p}_{m}^{[v]})]\in\mathbb{C}^{(N+1)\times K}$ denotes the channel matrix between the $m$-th FCA and the users, and $\mathbf{n}_{m}^{[v]}\in\mathbb{C}^{1\times\tau}$ is the received Gaussian noise vector.
@@
In the centralized channel estimation scheme, for each pilot block $v$, every FCA sends its received signal $\mathbf{y}_{m}^{[v]}$ to the CPU. The CPU stacks the received pilots across the $M$ antennas as
@@
Moreover, according to (10), the effective channel from user $k$ to FCA $m$ at signal block $v$ can be written as
@@
The main steps of the proposed centralized channel estimation procedure are summarized in Algorithm 2. The computational complexity of Algorithm 2 is dominated by the OMP-based sparse recovery and scales as $\mathcal{O}\!\big(\sum_{k=1}^{K}L_{k}\,MV\,G\big)$, while constructing the dictionary incurs an additional $\mathcal{O}(MVGN)$ operations. Therefore, the overall complexity is $\mathcal{O}\!\big(\sum_{k=1}^{K}L_{k}\,MV\,G+MVGN\big)$. In terms of communication overhead, Algorithm 2 requires each FCA to upload its length-$\tau$ received pilot vectors $\{\mathbf{y}_{m}^{[v]}\}_{v=1}^{V}$ to the CPU over $V$ pilot blocks, resulting in a communication overhead on the order of $\mathcal{O}(MV\tau)$. The centralized channel estimation scheme provides high estimation accuracy since the CPU jointly processes the pilot-correlated observations stacked over all FCAs and all pilot blocks.
Algorithm 2 Centralized channel estimation for FCA system
@@
3:  Each FCA $m$ applies $\mathbf{p}_{m}^{[v]}$ and receives $\mathbf{y}_{m}^{[v]}$ in (73).
4:  Each FCA $m$ uploads $\mathbf{y}_{m}^{[v]}$ to the CPU.
@@ IV-B Distributed Channel Estimation
The centralized channel estimation approach suffers from prohibitively high computational complexity, processing cost, and latency at the CPU, especially when the BS is equipped with an extremely large FCA array with a large number of active antennas $M$ and a vast number of coupler candidate positions. We aim to design a distributed channel estimation scheme for the FCA array that achieves accuracy close to the centralized baseline, while reducing the CPU computational burden.
@@ IV-B1 Processing at local nodes
Similar to the centralized channel estimation step, we discretize the azimuth domain into $G\gg L_{k}$ grid points $\{\phi_{1},\ldots,\phi_{G}\}$ and define the local dictionary at FCA $m$ as
@@
where $\mathbf{x}_{m,k}\in\mathbb{C}^{G\times 1}$ is sparse and its support encodes the azimuth angles of user $k$ on the grid. The support is common across $m$ since these angles are global parameters shared by all FCAs. Moreover, $\mathbf{n}_{m,k}\!\sim\!\mathcal{CN}(\mathbf{0},\sigma_{m,\mathrm{eff}}^{2}\mathbf{I}_{V})$ denotes the effective noise vector after pilot correlation and stacking over the $V$ pilot blocks, with $\sigma_{m,\mathrm{eff}}^{2}=\sigma_{m}^{2}/\tau$.
To convey support information with low overhead, FCA $m$ computes a matched-filter proxy $\mathbf{c}_{m,k}=\mathbf{A}_{m}^{\mathsf{H}}\widehat{\mathbf{y}}_{m,k}$ and its normalized energy $\rho_{m,k}(j)=\frac{|[\mathbf{c}_{m,k}]_{j}|^{2}}{\|\mathbf{a}_{m}(\phi_{j})\|_{2}^{2}+\epsilon_{\mathrm{n}}}$, $j=1,\ldots,G$, where $\epsilon_{\mathrm{n}}>0$ avoids numerical issues when $\|\mathbf{a}_{m}(\phi_{j})\|_{2}^{2}$ is small. It then selects a candidate support set via noise-calibrated hard thresholding [6]:
@@
where $\eta>0$ controls the pruning aggressiveness. Finally, FCA $m$ uploads only the index-value pairs $\mathcal{U}_{m,k}=\{(j,\rho_{m,k}(j))\}_{j\in\mathcal{I}_{m,k}}$ to the CPU. Since the azimuth domain is sparse, the uploaded list is typically low-dimensional, i.e., $|\mathcal{I}_{m,k}|\ll G$ in typical multipath channels.
@@ IV-B2 Processing at central node
Since azimuth angles $\{\phi_{k,\ell}\}_{\ell=1}^{L_{k}}$ are path-level geometric parameters and thus common to all FCAs, they are global parameters shared across different FCAs. Local index set $\{\mathcal{I}_{m,k}\}$ provides complementary evidence of this common support on the grid. Accordingly, the CPU forms an aggregated grid score for each $j\in\{1,\ldots,G\}$ as
@@ IV-B3 Post-processing at local nodes
The CPU broadcasts the estimated support set, $\widehat{\Omega}_{k}$, to all FCAs. Given $\widehat{\Omega}_{k}$, FCA $m$ restricts its local dictionary to the columns indexed by $\widehat{\Omega}_{k}$, namely,
@@
where $\bm{\alpha}_{k}\in\mathbb{C}^{L_{k}\times 1}$ collects the path gains aligned with the ordering of indices in $\widehat{\Omega}_{k}$. The reduced model allows each FCA to contribute to estimating the path gains using only a few coefficients, instead of processing the full $G$-dimensional sparse vector.
Algorithm 3 Distributed channel estimation for FCA systems
@@
4:  For each FCA $m$ in parallel, apply $\mathbf{p}_{m}^{[v]}$ and receive $\mathbf{y}_{m}^{[v]}$.
5:  For each FCA $m$ in parallel, compute $\widehat{g}_{m,k}^{[v]}$ for all $k$ via (87).
@@
7:for each FCA $m$ in parallel and each user $k$ do
@@
18:for each FCA $m$ in parallel and each user $k$ do
@@
To avoid sending the $V$-dimensional observation $\widehat{\mathbf{y}}_{m,k}$ from each FCA, FCA $m$ computes and forwards the following sufficient statistics for LS estimation to the CPU,
@@ IV-B4 Post-processing at central node
After receiving LS sufficient statistics $\{\mathbf{R}_{m,k},\mathbf{q}_{m,k}\}_{m=1}^{M}$ from all FCAs, the CPU aggregates them as
@@
After obtaining $\{\widehat{\Omega}_{k},\widehat{\bm{\alpha}}_{k}\}_{k=1}^{K}$ from the CPU, FCA $m$ reconstructs its effective scalar channel under any coupler position $\mathbf{p}_{m}$ as in (86). The reconstructed channel state information (CSI) is utilized for the subsequent coupler position optimization and digital precoding procedures described in Section III.
@@ V Simulation Results
Figure 3: Channel power gain (dB) between the FCA array and the user with $N=1$, $K=1$, $M=5$, $P_{\max}=30$ dBm, and $L_{k}=15$.
In this section, numerical results are presented to evaluate the performance of the proposed distributed FCA system design. The coupler movement region associated with each active antenna is modeled as a square with side length $A$. In the simulations, both the active antennas and the passive couplers are modeled as thin straight wire dipoles. The mutual impedance matrix is computed from closed-form expressions according to [1]. The active antennas are placed with an inter element spacing of $d_{y}=2.2\lambda$. Each coupler is allowed to move within a square region of side length $A=2\lambda$, subject to a minimum element spacing constraint of $d_{\min}=0.15\lambda$. The carrier frequency is set to $f_{c}=7$ GHz. The load-impedance matrix $\mathbf{X}$ is diagonal with identical entries $0.05+j50~\Omega$.
Figure 4: Achievable rate versus maximum transmit power for different schemes with $N=2$, $K=3$, $M=15$, and $L_{k}=15$.
@@ V-A Performance of Coupler Position Optimization
To evaluate the performance of coupler position optimization, the proposed FCA scheme is compared with the following baseline schemes.
@@
  • Fixed-position coupler: The couplers are fixed at uniformly spaced positions with an inter-element spacing of $0.4\lambda$, and the load impedance matrix $\mathbf{X}$ has constant diagonal entries of $0.05+j50~\Omega$.
  • Electronically steerable parasitic array radiator (ESPAR)[10]: Closed-form optimization of load impedance matrix $\mathbf{X}$ is used to maximize the sum rate while keeping the coupler positions fixed at the same positions as in the fixed-position coupler case.
  • Fully active array: All $MN+M$ ports are configured as active antennas at fixed positions with half-wavelength spacing.
@@
Fig. 4 plots the achievable sum rate versus maximum transmit power $P_{\max}$ for the proposed FCA array and several baseline schemes. As $P_{\max}$ increases, all schemes achieve higher rates due to the improved effective signal-to-noise ratio (SNR). It is observed that the proposed FCA array closely approaches the fully active array with $MN+M$ RF chains while using only $M$ active antennas. This gain is achieved because coupler position optimization provides substantial mechanical beamforming and fading mitigation gains in rich multipath channels while maintaining a limited RF chain cost. In contrast, the ESPAR scheme optimizes only the load impedances with fixed coupler locations and therefore cannot fully exploit the spatial degree of freedom (DoF) offered by position reconfiguration. Moreover, the fixed position coupler and the active antenna array exhibit lower rates due to their fixed geometries and the lack of a reconfigurable aperture.
Figure 5: Achievable rate versus number of users with $N=2$, $M=10$, $P_{\max}=30$ dBm, and $L_{k}=15$.
In Fig. 5, we plot the achievable sum rate versus the number of users $K$ for the proposed FCA array and benchmark schemes. As $K$ increases, the sum rate improves for all schemes due to the increased spatial multiplexing gain, while the growth gradually slows down since stronger multiuser interference must be mitigated under a fixed transmit power budget. Compared with the ESPAR and fixed position coupler schemes, the proposed design consistently achieves a higher sum rate. This performance advantage becomes more pronounced at moderate and large values of $K$, where MMSE based multiuser precoding is more sensitive to the conditioning of the effective channel matrix and thus benefits more from geometry reconfiguration, whereas the benchmark schemes with fixed geometries cannot adapt to user induced interference coupling.